OpenStudy (anonymous):

Jackie is analyzing a quadratic function f(x) and a linear function g(x). Will they intersect?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

a quadratic function f(x), can be this: \[f\left( x \right) = a{x^2} + bx + c\] whereas a linear function, g(x), can be this: \[g\left( x \right) = kx + h\]

OpenStudy (michele_laino):

now, the intersections between those functions, are given by the solutions of the subsequent algebraic system: \[\left\{ \begin{gathered} f\left( x \right) = a{x^2} + bx + c \hfill \\ g\left( x \right) = kx + h \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

so, we can have the subsequent cases: 1) two distinct solutions, which means, taht we have two distinct point of intersections: